# ChartMimic: Evaluating LMM’s Cross-Modal Reasoning Capability via Chart-to-Code Generation

### v1

@@ Appendix E Correlation with Human Evaluation
In order to evaluate the reliability of the proposed automatic metrics, we calculate their correlation with human evaluations. Specifically, a selection of $300$ test examples from the Direct Mimic task is utilized. Subsequently, we gather the results generated by GPT-4V using four different prompting methods (Sec. 4.2) on these test examples. This process yields $1,200$ figures to be assessed.
@@
Appendix F Cases of Error Analysis
@@
Appendix G Ethics and Societal Impact
We envision ChartMimic as a comprehensive suite of benchmarks designed to guide researchers in assessing the capabilities of their LMMs. By providing an extensive evaluation framework, ChartMimic aims to foster the development of foundation models for the community, offering insights into various aspects of LMMs, such as visual understanding, code generation, and cross-modal reasoning.
When constructing prototype charts, we use figures from papers on arXiv, which are licensed under CC BY 4.0. During the manual annotation of the code, we implement strict ethical controls to ensure that the information conveyed in the code does not contain any biased content and does not infringe on any private data.

### v2

@@ Appendix G Correlation with Human Evaluation
In order to evaluate the reliability of the proposed automatic metrics, we calculate their correlation with human evaluations. Specifically, a selection of $300$ test examples from the Direct Mimic task is utilized. Subsequently, we gather the results generated by GPT-4o using four different prompting methods (Sec. 4.2) on these test examples. This process yields $1,200$ figures to be assessed.
@@
Appendix H Correlation with Chart Understanding and Code Generation
Figure 49: Performance correlation of ChartMimic with benchmarks assessing chart understanding and code generation capabilities.
In an effort to shed light on the factors enhancing performance on ChartMimic, we explore the performance correlation between ChartMimic and existing benchmarks that evaluate chart understanding and code generation capabilities. Specifically, we select ChartQA (Masry et al., 2022) as the benchmark for chart understanding and HumanEval (Chen et al., 2021) as the benchmark for code generation. We then calculate two Pearson correlation coefficients to quantify these relationships: The correlation coefficient between the performance of LMMs in the Direct Mimic task and their performance in the ChartQA task, denoted as $r_{chart}$, and the correlation coefficient between the performance of LMMs in the Direct Mimic task and the performance of their corresponding LLMs in the HumanEval task, denoted as $r_{code}$
As depicted in Fig. 49, the calculated values are $r_{chart}$ = 0.6877 and $r_{code}$ = 0.8714. These results indicate that both chart understanding and code generation abilities influence the performance in the Direct Mimic task, with code generation having a more significant impact. This implies that for future model development aimed at enhancing multimodal code generation capabilities, it is crucial to focus on the foundational code generation abilities of LMMs.
Appendix I Cases of Error Analysis
@@
Appendix J Error Analysis of Open-Weight Models
Figure 51: Error analysis of InternVL2-Llama3-76B across four error types on the Direct Mimic task. The number in brackets indicates the count of error case.
Figure 52: Error analysis of InternVL2-26B across four error types on the Direct Mimic task. The number in brackets indicates the count of error case.
Figure 53: Error analysis of DeepSeek-VL-7B across four error types on the Direct Mimic task. The number in brackets indicates the count of error case.
We conduct detailed error analysis on representative open-weight models of different scales, sampling 100 cases for each model. For large-scale models such as InternVL2-Llama3-76B and InternVL2-26B (Fig. 51Fig. 52), the error patterns align closely with those observed in GPT-4o. Specifically, Code-related Errors predominantly arise from dimensional inconsistencies, while Text-related Errors primarily manifest as missing elements. Type-related Errors mainly comprise missing components cases, and Color recognition demonstrates the capability to identify similar, if not identical, chromatic properties. In contrast, smaller-scale models like DeepSeek-VL-7B (Fig. 53) exhibit markedly different error distributions. Code-related issues show a significant increase in Access and Parameter Errors (approximately 53.7%), while Text-related errors demonstrate a higher prevalence of missing elements (approximately 66.7%). Type-related cases reveal a substantial increase in missing errors (approximately 78.4%). Moreover, the degradation in color recognition is more severe, with the model failing to identify even similar colors. These findings indicate that while larger models encounter challenges similar to GPT-4o, smaller models exhibit more fundamental difficulties in both code generation and visual comprehension.
Appendix K Ethics, Societal Impact and Scalability of ChartMimic
ChartMimic is designed as a comprehensive evaluation framework for advancing LMMs, with a strong emphasis on ethical considerations, societal impact and scalability. Its primary objective is to provide researchers with a rigorous, ethically sound tool for assessing LMMs’ capabilities across critical domains, including visual understanding, code generation and cross-modal reasoning.
Ethical Considerations.
The ethical integrity of ChartMimic is foundational to its design and implementation. Our dataset is derived from scientific domains, specifically using figures from arXiv papers distributed under the CC BY 4.0 license. This focus on scientific content significantly mitigates potential ethical concerns related to bias, representation, or sensitive personal information that often arise in datasets derived from social media or general web content. The scientific nature of our data ensures: Minimal risk of perpetuating societal biases or stereotypes; Absence of personally identifiable information; Content that is generally neutral and objective, focusing on factual scientific representations. Our manual annotation process for code generation incorporates stringent ethical controls. Annotators are trained to ensure that the information conveyed in the code does not contain any biased content and strictly adheres to the factual representation of the scientific charts. This process further reinforces the ethical robustness of our dataset.
Societal Impact.
By providing a benchmark rooted in scientific data, ChartMimic contributes to the development of LMMs that can better understand and interact with complex, data-driven visualizations. This capability has far-reaching positive implications for scientific communication, data analysis, and knowledge dissemination across various fields.
Scalability.
The architecture of ChartMimic is inherently extensible, featuring a modular codebase that facilitates the seamless integration of additional chart types and evaluation metrics. Our data collection methodology leverages the continuous update cycle of arXiv, enabling sustainable expansion of the dataset while maintaining its focus on ethically sound, scientific content. The evaluation framework also demonstrates notable scalability: The concept of low-level metric can be readily adapted to other programming environments such as JavaScript and R. The high-level metric incorporates a meta-evaluation approach using LMMs, which allows for sustainable alignment with human preferences as stronger models emerge.